American Corporation has two equal shareholders, Mr. Freedom and Brave Inc. In addition to their investments in American stock, both shareholders have made substantial loans to American. During the current year, American paid $140,000 interest each to Mr. Freedom and Brave Inc. Assume that American and Brave have 21 percent tax rates, and Mr. Freedom’s marginal tax rate on ordinary income is 37 percent.
A)Calculate American’s tax savings from deduction of these interest payments and their after-tax cost.
B)Calculate Brave’s tax cost and after-tax earnings from its receipt of interest income from American.
C)Calculate Mr. Freedom’s tax cost and after-tax earnings from his receipt of interest income from American.
E)Recalculate Brave’s tax cost and after-tax earnings assuming its receipt of interest from American is treated as a constructive dividend.
F)Recalculate Mr. Freedom’s tax cost and after-tax earnings assuming his receipt of interest from American is treated as a constructive dividend.
**Please help me with part A and part F please as those 2 parts(4 numbers) are incorrect:
a.
Tax savings
$29,400 selected answer incorrect

After-tax cost of interest
$110,600 selected answer incorrect
b.
Tax cost
$29,400 selected answer correct

After-tax earnings
$110,600 selected answer correct
c.
Tax cost
$51,800 selected answer correct

After-tax earnings
$88,200 selected answer correct
e.
Tax cost
$10,290 selected answer correct

After-tax earnings
$129,710 selected answer correct
f.
Tax cost
$21,000 selected answer incorrect

After-tax earnings
$119,000 selected answer incorrect
Answer #1

(a) American’s tax savings from deduction of these interest payments and their after-tax cost.

Particulars Freedom Brave
Income from interest 140000 140000
Tax@ 21% 29400 29400
Marginal tax @37% 51800 -
Income after tax 58800 110600

American Tax saving due to interest= 29400+29400= 58800

(f) Calculation of Mr. Freedom’s tax cost and after-tax earnings assuming his receipt of interest from American is treated as a constructive dividend

Tax cost= 29400+51800= 81200

After tax earning= 140000-29400-51800= 58800

Note- American Inc can not claim corporate dividend as a business expense. and it shall also be liable for tax. And such dividend has no impact on the tax by freedom.